OpenStudy (anonymous):

Which value(s) must be excluded as a value of x in the polynomial? (x - 2 / (x2 - 25)

OpenStudy (anonymous):

so division by 0 is not alowed right ? so you have to find for which values of x you have division by 0 means to solve \[x^{2}-25=0\] can u do that ?

OpenStudy (anonymous):

lol no im stupid :/

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

nvm i basically forgot everything i learned this school year -,-

OpenStudy (anonymous):

if u dont know something doesnt mean ur stupid cmon xD so we have \[x^{2}-25=0\] if we add 25 to boths side of the equation we would have \[x^{2}-25+25=0+25\] which is \[x^{2}=25\] now we have square root of both side and we have \[\sqrt{x^{2}}=\sqrt{25}\]

OpenStudy (anonymous):

aw thank you lol

OpenStudy (anonymous):

do you know what values of x are excluded now ?

OpenStudy (anonymous):

if i say no will you hate me lol?

OpenStudy (anonymous):

no i will keep explaining until we get the answer :D

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

\[\sqrt{x^{2}}=\sqrt{25}\] ok so we have that , you know that square root of \[\sqrt{x^{2}} =x\] and \[\sqrt{25}=5\] so we get x=5 BUT we need to include x=-5 because \[(-5)^{2}=25\] so our final answer is x=5 and x=-5 should be excluded

OpenStudy (anonymous):

so the final answer is x=5 ?

OpenStudy (anonymous):

and x=-5

OpenStudy (anonymous):

both are ur answer

OpenStudy (anonymous):

okay tysm :)!

OpenStudy (anonymous):

no problem